# Evidence of 1:1 slope between rocky Super-Earths and their host stars

### v1

@@ 4.2 Multi-Planet systems
4.3 Role of galactic evolution
@@
Galactic Chemical Evolution (GCE) governs the metal content we observe in the stellar photospheres. In particular, the abundance ratios of elements (Mg, Si, Ca, Ti) to iron ([/Fe]) is simply put, a direct measurement of the different contributions of core-collapse supernovae (which favors the production of elements Timmes1995; Kobayashi2006) versus the contribution of thermonuclear explosions (type Ia supernovae - which contribute more significantly to the production of iron peak elements - Cr, Mn, Fe, Co, Ni, Cu, Zn Kobayashi2009; Kobayashi2020). The observed decrease in [/Fe] as stellar metallicity ([Fe/H]) increases is a direct result of the time-delay difference between core collapse and type Ia supernovae. Therefore, in a nutshell, most elements are produced by massive stars (MM⊙) within 20-30 Myrs, while Fe is produced primarily on a much longer timescale (up to 1 Gyr, see Kobayashi:2020ApJ...900..179K, for a GCE review). From the perspective of GCE, [/Fe] can, therefore, be a proxy of how late or early a star has formed. It is important to note, that this is a simplified view that does not account for the complex mixing processes of the interstellar medium (ISM) or the time delay required for the ISM to homogenize, prior to formation of new generation of stars, nor the specific details of the chemical evolution of the galactic substructure where the star was formed (i.e. thin disk, thick disk, bulge, halo).
@@
In these simple terms though and referring to trends applicable for a large population of stars, most stars formed in the thick disk are -rich and possibly older, compared to thin disk stars, which mostly have Sun-like [/Fe] ratios (Gilmore1983; Matteucci:1989epg..conf..297M; Fuhrmann1998; Kobayashi:2020ApJ...900..179K). The expectation then is that the Fe-MFpl of exoplanets and stellar Fe-MFst values, which is directly proportional to [/Fe] ratios, could have a correlation with stellar age.
@@
We investigate this and find that in our sample neither the Fe-MFst nor individual abundances of stars are correlated with age. In contrast for planets, however, we find a negative slope between Fe-MFpl and age of (see Figure 10), in line with iron enrichment in the ISM over time. 555This relationship holds irrespective of whether we use our weighted sample or include the two RTR exoplanets with equal weights or exclude them altogether (, see table 3). The former outcome is unsurprising, given that the stars in our sample span a limited range in both [Fe/H] and [/Fe] values, consistent with their common Galactic substructure origin. Consequently, the absence of a correlation likely reflects limited precision as the intrinsic stellar differences are expected to be small (Lorenzo2018). Furthermore, the stellar ages in our sample are derived from isochrone analysis alone, which does not account for systematics and thus our stellar ages require benchmarking to other methodologies for robustness. For example, using chemical clock analysis (Shejeelammal2024), we find that 55 Cnc has a drastically different stellar age of Gyr in comparison to our isochrone result of 12.30 Gyr. These factors introduce intrinsic challenges of measuring stellar ages, we refer the reader to Reggiani2022; Reggiani2024 for a detailed description of our methodology and soderblom2010 for a comprehensive review of the difficulties in estimating the ages of field stars.
@@
For the latter outcome on planets, there is possibly also a source of strong bias. Radial-velocity jitter and hence the observed uncertainty, scales with stellar age (Brems2019), with younger stars exhibiting higher jitter. Meaning, there may exists a strong bias toward planets with lower densities and lower Fe-MF, since for a given radius () planets that straddle the RTR boundary (CMF) are only detectable around older ( 5 Gyr) host stars (see paperI for details). Thus, these linear fits obtained with a few dozen planetary systems are most likely an overestimate of the correlation strength. Instead of stellar ages, hence, we compare the host star [] as well as [] abundances to the Fe-MFpl, which we discover to not have as robust of a linear relationship. We find that Fe-MFpl exhibits a negative correlation with [/Fe] (), while its relationship with [/H] yields a positive slope (). In both cases, however, no correlation () is consistent within . These findings raise interesting questions about whether planetary composition changes with galactic evolution and if older stars are more likely to host certain types of planets. Necessary improvements for this test require accurate and robust stellar ages or kinematics to determine thin/thick disk provenance (outside the scope of our paper) as well as a sample size that includes stars at a wider range of abundances ([Fe/H] or [Fe/H], which will depend on upcoming planetary discovery missions).
5 Conclusions
@@
We uncover a near 1:1 relationship between planets and stars in Fe-MF space, Fe-MF Fe-MF. This result suggests a primordial origin composition link when looking at the ensemble of planets. We find this relationship is consistent with data from Brinkman2024 when the planets that straddle the volatile-rocky composition boundary are properly treated, while it is in disagreement with data from Adibekyan2021/Adibekyan2024.
@@
| Name | M† (M⊕) | R† (R⊕) | CMF‡ | CMF | Fe-MF | (w) | (w) | (w) | |
| 55 Cnc e | - | - | 0.68 | 11.7 | - | - | - | - | - |
| HD 136352 b | 4.96 | 1.61 | 0.83 | 40.6 | 10.0 | 13.6 | 0.58 | 0.78 | 0.35 |
| HD 15337 b | 7.73 | 1.72 | 1.00 | 41.2 | 28.8 | 28.8 | 1.52 | 1.97 | 0.93 |
| HD 213885 b | 8.18 | 1.74 | 1.01 | 35.4 | 30.8 | 30.9 | 1.88 | 1.81 | 0.98 |
| HD 219134 b | 4.72 | 1.56 | 0.91 | 15.0 | 18.3 | 20.4 | 0.95 | 1.24 | 0.58 |
| HD 219134 c | 4.35 | 1.48 | 1.02 | 24.9 | 32.6 | 32.1 | 1.76 | 2.29 | 1.07 |
| HD 80653 b | 5.78 | 1.60 | 0.98 | 28.8 | 29.1 | 29.2 | 1.82 | 1.59 | 0.91 |
| K2-106 b | 8.51 | 1.78 | 0.88 | 63.5 | 25.0 | 25.9 | 1.46 | 1.37 | 0.75 |
| K2-131 b | 8.54 | 1.71 | 0.99 | 27.8 | 39.3 | 37.7 | 2.57 | 2.41 | 1.32 |
| K2-141 b | 5.36 | 1.59 | 0.96 | 39.3 | 23.7 | 25.1 | 1.81 | 1.06 | 0.73 |
| K2-229 b | 2.74 | 1.26 | 1.13 | 60.8 | 44.9 | 41.2 | 3.26 | 2.80 | 1.63 |
| K2-265 b | 6.70 | 1.62 | 1.09 | 31.4 | 38.2 | 36.1 | 2.64 | 2.11 | 1.28 |
| K2-291 b | 6.38 | 1.57 | 1.14 | 44.2 | 44.6 | 42.2 | 2.97 | 2.94 | 1.57 |
| K2-38 b | 7.50 | 1.62 | 1.21 | 55.5 | 48.6 | 44.6 | 3.29 | 3.62 | 1.86 |
| Kepler-107 c | 9.09 | 1.59 | 1.57 | 66.2 | 72.7 | 64.7 | 7.84 | 7.63 | 4.19 |
| Kepler-10 b | 3.37 | 1.47 | 0.82 | 13.1 | 6.2 | 9.9 | 0.39 | 0.58 | 0.24 |
| Kepler-20 b | 9.41 | 1.77 | 1.07 | 26.3 | 37.2 | 36.2 | 2.34 | 2.38 | 1.26 |
| Kepler-36 b | 3.67 | 1.51 | 0.77 | 28.1 | 5.9 | 10.0 | 0.44 | 0.42 | 0.22 |
| Kepler-78 b | 1.82 | 1.14 | 0.97 | 37.9 | 40.4 | 38.4 | 2.37 | 2.77 | 1.37 |
| Kepler-93 b | 4.52 | 1.51 | 1.01 | 26.7 | 29.5 | 30.0 | 1.47 | 2.25 | 0.94 |
| WASP-47 e | 7.69 | 1.83 | 0.74 | 7.4 | 4.2 | 7.2 | 0.31 | 0.29 | 0.16 |
Posterior results for Mass and Radius data, may be different from the observational data owing to sampling with MCMC as some distributions are outside the rocky region (RTR).
Planetary CMF data from Plotnykov2020.
| Name | M (M⊙) | R (R⊙) | [Fe/H] (dex) | [Mg/H] (dex) | [Si/H] (dex) | Fe-MFst | Age (Gyr) |
| 55 Cnc | 0.92 | 1.00 | 0.42 | 0.53 | 0.48 | 26.8 | 12.30 |
| HD 136352 | 0.88 | 1.03 | -0.22 | -0.02 | -0.18 | 25.3 | 10.71 |
| HD 15337 | 0.90 | 0.85 | 0.23 | 0.38 | 0.24 | 26.8 | 5.10 |
| HD 213885 | 0.96 | 1.10 | 0.05 | 0.05 | 0.04 | 30.9 | 7.65 |
| HD 219134 | 0.80 | 0.76 | 0.21 | 0.38 | 0.24 | 26.2 | 6.00 |
| HD 219134 | 0.80 | 0.76 | 0.21 | 0.38 | 0.24 | 26.2 | 6.00 |
| HD 80653 | 1.19 | 1.22 | 0.35 | 0.29 | 0.32 | 32.6 | 2.61 |
| K2-106 | 0.93 | 0.99 | 0.10 | 0.09 | 0.10 | 30.9 | 9.48 |
| K2-131 | 0.83 | 0.76 | -0.09 | -0.04 | -0.03 | 28.1 | 2.26 |
| K2-141 | 0.79 | 0.72 | 0.28 | 0.16 | 0.38 | 29.9 | 1.36 |
| K2-229 | 0.89 | 0.80 | 0.12 | 0.20 | 0.34 | 23.7 | 1.04 |
| K2-265 | 0.95 | 0.93 | 0.16 | 0.22 | 0.29 | 26.3 | 5.67 |
| K2-291 | 0.92 | 0.89 | 0.09 | 0.07 | 0.06 | 32.2 | 4.39 |
| K2-38 | 1.02 | 1.15 | 0.22 | 0.30 | 0.23 | 28.5 | 8.23 |
| Kepler-107 | 1.25 | 1.44 | 0.31 | 0.35 | 0.32 | 29.7 | 3.72 |
| Kepler-10 | 0.92 | 1.08 | -0.08 | 0.12 | -0.11 | 26.6 | 10.19 |
| Kepler-20 | 0.89 | 0.91 | 0.05 | 0.03 | 0.00 | 32.5 | 7.93 |
| Kepler-36 | 1.09 | 1.70 | -0.16 | -0.15 | -0.17 | 31.3 | 6.60 |
| Kepler-78 | 0.82 | 0.75 | 0.03 | 0.15 | 0.06 | 27.2 | 1.50 |
| Kepler-93 | 0.88 | 0.94 | -0.09 | 0.03 | -0.18 | 29.8 | 8.48 |
| WASP-47 | 1.03 | 1.16 | 0.41 | 0.48 | 0.45 | 27.8 | 8.58 |
| With weights () | Equal weights () | No RTR planets () |
| Fe-MFpl = 0.97 Fe-MFst 0.03 | 1.83 Fe-MFst 0.29 | 1.29 Fe-MFst 0.10 |
| Fe-MFpl = 0.24 [Fe/H] 0.23 | 0.11 [Fe/H] 0.20 | 0.37 [Fe/H] 0.23 |
| Fe-MFpl = 0.13 [Mg/H] 0.23 | -0.01 [Mg/H] 0.22 | 0.20 [Mg/H] 0.23 |
| Fe-MFpl = 0.19 [Si/H] 0.23 | 0.07 [Si/H] 0.21 | 0.29 [Si/H] 0.23 |
| Fe-MFpl = -0.46 [] 0.24 | -0.17 [] 0.28 | -0.65 [] 0.25 |
| Fe-MFpl = 0.15 0.24 | 0.03 0.21 | 0.24 0.25 |
| Fe-MFpl = -0.02 Age 0.38 | -0.02 Age 0.38 | -0.02 Age 0.40 |
| 0.97 Fe-MFst 0.65 | 2.59 Fe-MFst 0.17 | 1.25 Fe-MFst 0.60 |
| 0.21 [Fe/H] 0.91 | -0.26 [Fe/H] 0.90 | 0.32 [Fe/H] 0.93 |
| 0.13 [Mg/H] 0.91 | -0.26 [Mg/H] 0.92 | 0.22 [Mg/H] 0.92 |
| 0.15 [Si/H] 0.92 | -0.22 [Si/H] 0.90 | 0.24 [Si/H] 0.93 |
| -0.45 [] 0.92 | -1.25 [] 0.87 | -0.54 [] 0.94 |
| -0.04 [/H] + 0.94 | 0.39 [/H] + 0.85 | -0.35 [/H] + 1.02 |
| -0.02 Age 1.04 | -0.03 Age 1.09 | -0.02 Age 1.06 |

### v2

@@ IV.2 Multi-Planet systems
IV.3 Role of galactic evolution
@@
Galactic Chemical Evolution (GCE) governs the metal content we observe in the stellar photospheres. In particular, the abundance ratios of elements (Mg, Si, Ca, Ti) to iron ([/Fe]) is simply put, a direct measurement of the different contributions of core-collapse supernovae (which favors the production of elements Timmes et al. 1995; Kobayashi et al. 2006) versus the contribution of thermonuclear explosions (type Ia supernovae - which contribute more significantly to the production of iron peak elements - Cr, Mn, Fe, Co, Ni, Cu, Zn Kobayashi and Nomoto 2009; Kobayashi et al. 2020b). The observed decrease in [/Fe] as stellar metallicity ([Fe/H]) increases is a direct result of the time-delay difference between core collapse and type Ia supernovae. Therefore, in a nutshell, most elements are produced by massive stars (MM⊙) within 20-30 Myrs, while Fe is produced primarily on a much longer timescale (up to 1 Gyr, see Kobayashi et al., 2020a, for a GCE review). From the perspective of GCE, [/Fe] can, therefore, be a proxy of how late or early a star has formed. It is important to note, that this is a simplified view that does not account for the complex mixing processes of the interstellar medium (ISM) or the time delay required for the ISM to homogenize, prior to formation of new generation of stars, nor the specific details of the chemical evolution of the galactic substructure where the star was formed (i.e. thin disk, thick disk, bulge, halo).
@@
In these simple terms though and referring to trends applicable for a large population of stars, most stars formed in the thick disk are -rich and possibly older, compared to thin disk stars, which mostly have Sun-like [/Fe] ratios (Gilmore and Reid, 1983; Matteucci, 1989; Fuhrmann, 1998; Kobayashi et al., 2020a). The expectation then is that the Fe-MFpl of exoplanets and stellar Fe-MFst values, which is directly proportional to [/Fe] ratios, could have a correlation with stellar age or be affected by GCE as shown and argued by Weeks et al. (2025); Behmard et al. (2025).
@@
We investigate this and find that in our sample neither the Fe-MFst nor individual abundances of stars are correlated with age. In contrast for planets, however, we find a negative slope between Fe-MFpl and age of (see Figure 10), in line with iron enrichment in the ISM over time and in agreement with Weeks et al. (2025) results. The former outcome is unsurprising, given that the stars in our sample span a limited range in both [Fe/H] and [/Fe] values, consistent with their common Galactic substructure origin. Consequently, the absence of a correlation likely reflects limited precision as the intrinsic stellar differences are expected to be small (Spina et al., 2018). Furthermore, the stellar ages in our sample are derived from isochrone analysis alone, which does not account for systematics and thus our stellar ages require benchmarking to other methodologies for robustness. For example, using chemical clock analysis (Shejeelammal et al., 2024), we find that 55 Cnc has a drastically different stellar age of Gyr in comparison to our isochrone result of 12.30 Gyr. These factors introduce intrinsic challenges of measuring stellar ages, we refer the reader to Reggiani et al. (2022, 2024) for a detailed description of our methodology and Soderblom (2010) for a comprehensive review of the difficulties in estimating the ages of field stars.
@@
For the latter outcome on planets, there is possibly also a source of strong bias. Radial-velocity jitter and hence the observed uncertainty, scales with stellar age (Brems et al., 2019), with younger stars exhibiting higher jitter. Meaning, there may exists a strong bias toward planets with lower densities and lower Fe-MF, since for a given radius () planets that straddle the RTR boundary (CMF) are only detectable around older ( 5 Gyr) host stars (see paper I for details). Thus, these linear fits obtained with a few dozen planetary systems are most likely an overestimate of the correlation strength. Instead of stellar ages, hence, we compare the host star [] as well as [] abundances to the Fe-MFpl, which we discover to not have as robust of a linear relationship. We find that Fe-MFpl exhibits a negative correlation with [/Fe] (), while its relationship with [/H] yields a positive slope (). In both cases, however, no correlation () is consistent within . These findings raise interesting questions about whether planetary composition changes with galactic evolution and if older stars are more likely to host certain types of planets. Necessary improvements for this test require accurate and robust stellar ages or kinematics to determine thin/thick disk provenance (outside the scope of our paper) as well as a sample size that includes stars at a wider range of abundances ([Fe/H] or [Fe/H], which will depend on upcoming planetary discovery missions).
V Conclusions
@@
We uncover a near 1:1 relationship between planets and stars in Fe-MF space, Fe-MF Fe-MF. This result suggests a primordial origin composition link when looking at the ensemble of planets. We find this relationship is consistent with data from B24 when the planets that straddle the volatile-rocky composition boundary are properly treated, while it is in disagreement with data from A21/A24.